OpenStudy (anonymous):

Solve the following: X=5y-3 X+y=9

OpenStudy (anonymous):

x=5y-3 x+y=9 5y-3)-y=9 6y-3=9 6y=12 y=2 x=5(2)-3 x=10-3 x=7

OpenStudy (lilg132):

x = 5y - 3 x + y = 9 -x + 5y = 3 x + y = 9 6y = 12 6y / 6 = 12 / 6 y = 2

OpenStudy (lilg132):

x = 5(2) - 3 x = 10 - 3 x = 7